OpenStudy (anonymous):

Find the radius, diameter and circumference of a circle with an area of 64 cm2. Write your answers as decimals rounded to the nearest hundredth.

OpenStudy (anonymous):

hey homi

OpenStudy (anonymous):

whaddup G :D

OpenStudy (anonymous):

good good wat do u need

OpenStudy (anonymous):

i need to Find the radius, diameter and circumference of a circle with an area of 64 cm2. Write your answers as decimals rounded to the nearest hundredth.

OpenStudy (anonymous):

i suck at math im mad slow in this subject soo yeahh haha

OpenStudy (anonymous):

i dont like circles but i will try to help you

OpenStudy (anonymous):

blahh me either -__-

OpenStudy (anonymous):

^n^

OpenStudy (anonymous):

ok so your area is ...

OpenStudy (anonymous):

64cm^2

OpenStudy (anonymous):

64 cm sqrd

OpenStudy (anonymous):

well cut that in half

OpenStudy (anonymous):

64/2

OpenStudy (anonymous):

ok 8cm^2

OpenStudy (anonymous):

32 sorry sorry

OpenStudy (anonymous):

lol i was thinking something else haha

OpenStudy (anonymous):

thats is ur diameter (i think)

OpenStudy (anonymous):

32/2

OpenStudy (unklerhaukus):

\[A=\pi r^2\] \[\sqrt{\frac{A}\pi}= r\] \[C=2\pi r\]\[r=d/2\] \[\sqrt{\frac{A}\pi}= r\]

OpenStudy (anonymous):

oh well there u have it i think i waz going in the wrong direction

OpenStudy (anonymous):

haha its all good xD. but i dont even have a calculator to do that -__-

OpenStudy (anonymous):

me niether

OpenStudy (anonymous):

no calculator required for this one

OpenStudy (anonymous):

like i said im dumb in math haha

OpenStudy (anonymous):

follow the formula that unklerhakus typed in

OpenStudy (anonymous):

oh